# thanks: Corresponding author

### v1

@@ III.3 Transfer functions
To determine the optical appearance of a BH surrounded by a thin disk at various inclination angles, we must establish the mapping between each point on the accretion disk and its corresponding position on the observer’s plane. Here, we introduce the transfer function to characterize the relationship between points on the observer’s plane and their originating points on the accretion disk [31]. Notably, since light rays from the counter-side and co-side semi-disks undergo different deflection angles and cannot be treated uniformly, we classify and discuss the transfer functions separately.
For the the counter-side semi-disk, its transfer function $r_{\rm m}(b)$ can be expressed as:
$r_{\rm m}(b)=\frac{1}{u\left[\frac{m\pi}{2}+\arctan{({\tan{\omega}\sin{\eta}})},b\right]},\qquad m=1,2,3,\cdots,$ (11)
while for the co-side semi-disk, the transfer function $R_{\rm m}(b)$ takes the form:
$R_{\rm m}(b)=\frac{1}{u\left[\frac{m\pi}{2}-\arctan{({\tan{\omega}\sin{\eta}})},b\right]},\qquad m=1,2,3,\cdots.$ (12)
Here, $m$ denotes the radial position at the $m$-th intersection point between the light ray and the accretion disk. Notably, for the co-side transfer function, photons with $m=2$ actually intersect the counter-side semi-disk, yet their contributions appear in the co-side image. Therefore, the co-side (or counter-side) transfer functions essentially characterize the intersection positions on the accretion disk of light rays received at each point on the observer’s plane. Specifically, these transfer functions represent the contributions from different disk radii $r$ to the image. We derive these transfer functions by numerically solving Eq. 7.
Similar to previous analyses, Figs. 6 and 7 respectively present the transfer functions for the counter-side semi-disk and the first three transfer functions for the co-side semi-disk, with fixed $\eta=\pi/2$ and $\omega=17\pi/36$. Here, black, orange, and red curves correspond to $m=1$, $m=2$, and $m=3$ cases. The results align with prior findings on orbit numbers: as inclination $\omega$ (or polar angle $\eta$) increases, the impact parameter ranges for the second and third transfer functions of the counter-side semi-disk shrink, while those of the co-side semi-disk expand. Moreover, the slope of the first transfer function grows progressively with increasing $\omega$ and $\eta$, accompanied by a reduction in its impact parameter range–indicating contraction of the direct emission region on the observer’s plane at higher viewing angles. Notably, when the inclination approaches $\pi/2$, the lensed emission region of the co-side semi-disk becomes dozens of times larger than its counter-side counterpart. This implies that for moderately sized accretion disks, the co-side emission will be predominantly sourced from the lensed emission region.
@@
IV Optical Appearance of Black Hole at Different Inclinations with Thin Disks
In the previous section, we established the correspondence between emission points on the accretion disk and their projected positions on the observer’s plane. Incorporating both the local emission intensity of the optically thin disk and gravitational redshift effects, the received intensity at frequency $\nu_{o}$ originating from radius $r$ with emitted frequency $\nu_{e}$ is given by [31]:
@@ IV Optical Appearance of Black Hole at Different Inclinations with Thin Disks
here $g$ denotes the redshift factor. The redshift factor can be calculated through $g=\frac{P_{\mu}u_{o}^{\mu}}{P_{\nu}u_{e}^{\nu}}$, where $P_{\mu}$ denotes the photon’s 4-momentum, while $u_{o}^{\mu}$ and $u_{e}^{\nu}$ correspond to the observer’s and the emitter’s 4-velocities [59], respectively. For a static observer at infinity, the 4-velocity is given by $u_{o}^{\mu}=(1,0,0,0)$, while for a static emitter, $u_{e}^{\nu}=(1/\sqrt{f(r)},0,0,0)$. The redshift factor is thus calculated as $g=\sqrt{f(r)}$. After integrating over the frequency $\nu_{e}$, we obtain:
@@
The observed brightness originates from the intersections between light rays and the accretion disk. Therefore, we sum contributions from all intersection points on both the counter-side and co-side semi-disks to obtain the final received intensity:
@@
Here, $I_{\rm obs1}(b)$ and $I_{\rm obs2}(b)$ represent the total received intensities from the counter-side and co-side semi-disks, respectively. As considered in [31], the $m=3$ transfer function has an extremely steep slope, contributing only to a drastically narrowed region of the image with negligible flux. We therefore exclude cases with $m>3$ in our analysis.
Thus, given a specific emission profile of the thin disk, we can derive the corresponding received intensity. We consider three commonly used emission models [31, 38]:
@@
where $I_{0}$ is the maximum emission intensity. Additionally, $r_{\rm h}$, $r_{\rm ph}$, and $r_{\rm isco}$ denote the event horizon radius, photon sphere radius, and innermost stable circular orbit (ISCO) radius of the Schwarzschild BH, respectively. Their specific calculations are omitted here; see [46, 52] for details. Figure 8 presents the emission intensity profiles for three distinct models: Model-I, Model-II, and Model-III, which correspond to emission profiles initiating their decay at $r_{\rm isco}$, $r_{\rm ph}$, and $r_{\rm h}$, respectively.
@@
Combining Eqs. (11), (12), and (15) we take Model-I as an example to demonstrate the received intensity on the observer’s plane at inclination angles of $\pi/4$ and $3617\pi/$ 36. Figures 9 and 10 show the intensity distributions for different polar angles ($\eta$) of the observer. From the figures, we observe that for the counter-side semi-disk, the number of intensity peaks in $I_{\rm obs1}$ shows no significant variation with polar angle $\eta$, while the distance between adjacent peaks displays only minor adjustments. Meanwhile, increasing the inclination angle reduces both the intensity at fixed $\eta$ and the peak separation distance. In contrast, the co-side intensity $I_{\rm obs2}$ exhibits more complex intensity profiles, developing additional emission peaks. For Model II and Model III, we briefly present the received intensity at select polar angles for inclination angles $\omega=\pi/4$ and $\omega=17\pi/36$ in Appendix A.
@@
Figure 11: Optical appearances of a Schwarzschild BH illuminated by three types of geometrically thin, optically thin accretion disks at varying inclination angles. Columns 1–3 display images for Model-I, Model-II, and Model-III at inclinations $\omega=\pi/36$, $18017\pi/$ 180, $\pi/4$, and $3617\pi/$ 36, respectively.
@@
V Optical Appearance of Black Hole at Different Inclinations with Thick Disks
We now examine the optical appearance of a Schwarzschild BH surrounded by optically thick versions of the three accretion disk models. Unlike the optically thin case, the optically thick disk obstructs photon propagation, preventing some rays from reaching the observer and resulting in darker images. Therefore, we need to explicitly determine the spatial location of the thick accretion disk.
Specifically, we define the radial extent of the optically thick disk as:
  • Model-I: $r_{\rm isco}\leq r\leq 12$
  • Model-II: $r_{\rm ph}\leq r\leq 12$
  • Model-III: $r_{\rm h}\leq r\leq 12$
@@ V Optical Appearance of Black Hole at Different Inclinations with Thick Disks
For Model-I, the transfer functions of the counter-side semi-disk (similarly for the co-side) in Fig. 7 reveal that within the impact parameter range corresponding to $r_{\rm isco}\leq r_{1}(b)\leq 12$, the $m=1$ component dominates the image contribution. Due to disk obscuration, the $m=2$ and $m=3$ transfer functions contribute negligibly to the intensity in this region. Conversely, outside this impact parameter range where obscuration is absent, if $r_{2}(b)$ and $r_{3}(b)$ satisfy $r_{\rm isco}\leq r\leq 12$, the lensed ($m=2$) and photon ring ($m=3$) emission regions will contribute significantly to the image intensity. Thus, similar to the previous procedure, we fix the inclination angle $\omega$ and analyze the transfer functions for different observer polar angles $\eta$. We then calculate the received intensity using Eq. (15), considering only those transfer functions that actually contribute to the observed image.
The same computational procedure applies to Model-II and Model-III. However, since these models adopt disk radii starting from $r_{\rm ph}$ and $r_{\rm h}$ respectively, the lensed ($m=2$) and photon ring ($m=3$) contributions become negligible in regions where $r<r_{\rm ph}$ or $r<r_{\rm h}$. This allows for certain simplifications in our analysis.
Here, we take $\omega=17\pi/36$ as an example to demonstrate the optical appearance of a Schwarzschild BH illuminated by thick accretion disks under the three models, with direct comparison to the thin-disk cases in Fig. 12. The top row of Fig. 12 displays the images for the three thin-disk models, while the bottom row shows their thick-disk counterparts. Two key effects are evident: (i) the bright rings disappear in the thick-disk images due to disk obscuration, and (ii) the overall intensity of the thick disks is approximately $20\%$$40\%$ lower than that of the thin disks.
@@ VI Conclusion
In this paper, we investigate the optical appearance of a Schwarzschild BH surrounded by three geometrically thin and optically thin accretion disk models as observed from different inclination angles. By projecting the observer’s plane (the $XOY^{\prime\prime}$ plane in Fig. 2) onto the BH system, we divide the accretion disk into two hemispheres: the co-side semi-disk (facing the observer) and the counter-side semi-disk (opposite the observer), and this classification is consistently used throughout our analysis.
@@
Following the methodology of [31], we categorize photon trajectories intersecting the disk once, twice, and three or more times as direct emission, lensed, and photon ring emissions, respectively. With this classification, each type of photon trajectory from both semi-disks produces distinct image components on the observer’s plane: the direct emission forms the primary image, while the lensed emission generates secondary images, and so forth. We then present the photon trajectories and orbit numbers for both semi-disks at fixed inclination $\omega$ and observer polar angle $\eta$. The results reveal striking differences between the two semi-disks: the lensed emission region is significantly larger on the co-side, and its size increases markedly with inclination, while the corresponding region on the counter-side shrinks under the same conditions.
@@
To establish the relationship between the accretion disk and points on the observer’s plane, we introduce transfer functions $r_{\rm m}(b)$ and $R_{\rm m}(b)$ for the counter-side and co-side semi-disks, respectively. These functions characterize the disk intersection positions of light rays received at each point on the observer’s plane. The variation of these transfer functions with inclination angle follows the same trend as the orbit numbers. Notably, when both the inclination $\omega$ and polar angle $\eta$ approach $\pi/2$, the lensed emission region of the co-side semi-disk dominates the image contribution.
@@
Subsequently, by specifying the emission profiles of the thin disk models and incorporating the transfer functions, we calculate the received intensity at different polar angles $\eta$ for a fixed inclination $\omega$. Our analysis reveals that the co-side semi-disk exhibits more complex peak structures in the received intensity across varying $\eta$, while the counter-side semi-disk maintains a nearly constant number of intensity peaks with only variations in their spacing. These results are then projected onto a two-dimensional plane to generate optical images for the three thin disk models at different inclinations, as shown in Fig. 11.
@@
Finally, we extend our discussion to optically thick versions of the three models, comparing their images with the thin disk cases using $\omega=17\pi/36$ as a representative example. The results demonstrate that the thick disk’s obscuration effect eliminates certain bright rings and reduces the overall intensity by approximately $20\%$$40\%$ compared to thin disks.
@@ Appendix A Received intensities of Model-II and Model-III
In this appendix, we present the received intensities of Model-II and Model-III in Figs. 1316.

### v2

@@ III.3 Transfer functions
To determine the optical appearance of a BH surrounded by a disk at various inclination angles, we must establish a map between each point on the accretion disk and its corresponding position on the observer’s plane. Indeed, the transfer function can be introduced to characterize the map [31]. Notably, since light rays from the counter-side and co-side semi-disks undergo different azimuthal angles and cannot be treated uniformly, one needs to consider the corresponding transfer functions separately. For the counter-side semi-disk, the transfer function $r_{\rm m}(b)\equiv r(\phi,b)=1/u(\phi,b)$, as a function of the inclination angle $\omega$ and the observer’s polar angle $\eta$, reads
$r_{\rm m}(b)=\frac{1}{u\left[\frac{2m-1}{2}\pi+\arctan{({\tan{\omega}\sin{\eta}})},b\right]},\;m=1,2,3,\cdots.$ (12)
Here, $m$ denotes the $m$-th intersection between the light ray and the accretion disk. Similarly, for the co-side semi-disk, the transfer function $R_{\rm m}(b)\equiv r(\phi,b)=1/u(\phi,b)$, which is also a function of $\omega$ and $\eta$, takes the form
$R_{\rm m}(b)=\frac{1}{u\left[\frac{2m-1}{2}\pi-\arctan{({\tan{\omega}\sin{\eta}})},b\right]},\;m=1,2,3,\cdots.$ (13)
It is worth to noting that the counter-side transfer function essentially characterizes the intersection positions on the entire accretion disk for light rays originating from each point in the observer’s upper half-plane, where these intersection points can also lie on the co-side disk. Conversely, the co-side transfer function describes the intersection positions on the entire accretion disk for light rays from each point in the observer’s lower half-plane. These transfer functions can be obtained by numerically solving Eq. (7).
Figures 7 and 8 respectively present the transfer functions for the counter-side semi-disk and the first three transfer functions for the co-side semi-disk, with fixed $\eta=\pi/2$ and $\omega=17\pi/36$. Here, black, orange, and red curves correspond to the cases of $m=1$, $m=2$, and $m=3$, respectively. The results align with prior findings on orbit numbers: as inclination $\omega$ (or polar angle $\eta$) increases, the impact parameter ranges for the second ($m=2$) and third ($m=3$) transfer functions of the counter-side semi-disk shrink, while those of the co-side semi-disk expand. Moreover, the slope of the first ($m=1$) transfer function grows progressively with increasing $\omega$ and $\eta$, accompanied by a reduction in its impact parameter range, which indicates a contraction of the direct emission region on the observer’s plane with a bigger $\omega$.
@@
IV Optical appearance of Schwarzschild black hole with a optically thin disk at different inclinations
In the previous section, we established the correspondence between emission points on the accretion disk and their projected positions on the observer’s plane by the transfer functions. In this section, we will consider both the local emission intensity of the optically thin disk and gravitational redshift effect, and obtain the received intensity at frequency $\nu_{o}$ originating from radius $r$ with emitted frequency $\nu_{e}$, which is given by [31]
@@ IV Optical appearance of Schwarzschild black hole with a optically thin disk at different inclinations
where $g$ denotes the redshift factor. The redshift factor can be calculated through $g=\frac{P_{\mu}u_{o}^{\mu}}{P_{\nu}u_{e}^{\nu}}$, where $P_{\mu}$ denotes the photon’s 4-momentum, while $u_{o}^{\mu}$ and $u_{e}^{\nu}$ correspond to the 4-velocities of the observer and the emitter, respectively [60]. For a static observer at infinity, its 4-velocity is given by $u_{o}^{\mu}=(1,0,0,0)$, while for a static emitter, its 4-velocity is $u_{e}^{\nu}=(1/\sqrt{f(r)},0,0,0)$. The redshift factor is thus calculated as $g=\sqrt{f(r)}$. After integrating over the frequency $\nu_{e}$, we obtain
@@
The observed intensity in the observer’s plane is determined by the total contribution of the emission intensities at the intersections between light rays and the accretion disk (neglecting absorption). Therefore, the final received intensities $I_{\rm obs1}(b)$ and $I_{\rm obs2}(b)$, as functions of $\omega$ and $\eta$, for the counter-side and co-side semi-disks read, respectively
@@
As pointed out in [31], the transfer function for $m=3$ has an extremely steep slope, contributing only to a drastically narrowed region of the image with negligible flux. We therefore exclude the cases of $m>3$ in our analysis.
Once the emission profile of the thin disk is specified, the corresponding received intensity can be obtained by Eq. (16). In this paper, we consider three commonly used emission models [31, 38]
@@
where $I_{0}$ is the maximum emission intensity, and the quantities $r_{\rm h}$, $r_{\rm ph}$, and $r_{\rm isco}$ denote the event horizon radius, photon sphere radius, and innermost stable circular orbit (ISCO) radius of the Schwarzschild BH, respectively. The values of these quantities are omitted here (see [46, 44, 52] for details). Figure 9 presents the emission intensity profiles for these three distinct models, which correspond to emission profiles initiating their decay at $r_{\rm isco}$, $r_{\rm ph}$, and $r_{\rm h}$, respectively.
@@
Combining Eqs. (12), (13), and (16), we calculate the received intensity on the observer’s plane for different $\eta$ at inclination angle $\omega$ with values $\pi/4$ and $3617\pi/$ 36 for the emission intensity characterized by the Model-I, and plotted the results in Fig. 10 and 11. From the figures, we observe that for the counter-side semi-disk, the number of intensity peaks in $I_{\rm obs1}$ shows no significant variation with polar angle $\eta$, while the distance between adjacent peaks displays only minor adjustments. Meanwhile, increasing the inclination angle reduces both the intensity at fixed $\eta$ and the peak separation distance. In contrast, the co-side intensity $I_{\rm obs2}$ exhibits more complex intensity profiles, developing additional emission peaks. For Model-II and Model-III, we briefly present the received intensities for select polar angles at inclination angles $\omega=\pi/4$ and $\omega=17\pi/36$ in Appendix A.
@@
Figure 12: Optical appearances of a Schwarzschild BH illuminated by three types of geometrically and optically thin accretion disks at varying inclination angles. Columns 1-3 display images for Model-I, Model-II, and Model-III at inclinations $\omega=\pi/36$, $18017\pi/$ 180, $\pi/4$, and $3617\pi/$ 36, respectively.
@@
V Optical appearance of Schwarzschild black hole with a optically thick disk at different inclinations
We now examine the optical appearance of a Schwarzschild BH surrounded by optically thick versions of the three accretion disk models. Unlike the optically thin case, the optically thick disk obstructs photon propagation, preventing some rays from reaching the observer and resulting in darker images. Therefore, to determine the final appearance, the spatial region of the thick accretion disk has to be specified. Based on the emission profiles of the three models, we define the disk’s inner boundaries at locations where the emission intensity drops to zero for the three emission models. Furthermore, due to the rapid emission decay in all models, we consider finite outer boundaries for these accretion disks. In what follows, we consider the regions of the optically thick disks with radius $r$ as
  • Model-I: $r_{\rm isco}\leq r\leq 12$.
  • Model-II: $r_{\rm ph}\leq r\leq 12$.
  • Model-III: $r_{\rm h}\leq r\leq 12$.
@@ V Optical appearance of Schwarzschild black hole with a optically thick disk at different inclinations
The analysis from previous sections remains applicable here. For Model-I, the impact parameter range in the observer’s plane, corresponding to the first transfer function $r_{1}(b)$ and $R_{1}(b)$ in $[r_{\rm isco},12]$, has no contributions from lensed or photon ring emission due to obstruction by the thick disk. However, outside this impact parameter range, the lensed and photon ring emission will contribute to the image intensity. Similar results hold for Model-II and Model-III. We therefore follow the earlier methodology, using Eq. (16) to calculate the received intensity for each transfer function.
In Fig. 13, we take $\omega=17\pi/36$ as an example to demonstrate the optical appearance of a Schwarzschild BH illuminated by thick accretion disks under the three models, with direct comparison to the thin-disk cases. The top row of Fig. 13 displays the images for the three thin-disk models, while the bottom row shows their thick-disk counterparts. Two key effects are evident: (i) the bright rings disappear in the thick-disk images due to disk obscuration, and (ii) the overall intensities of the thick disks is approximately $20\%$$40\%$ lower than those of the thin disks.
@@ VI Summary
In this paper, we investigate the optical appearance of a Schwarzschild BH surrounded by three geometrically thin accretion disk models as observed at different inclinations. Based on geometric configuration, we divided the accretion disk into two semi-disks: the co-side semi-disk (facing the observer) and the counter-side semi-disk (opposite the observer), and this classification is consistently used throughout our analysis.
@@
Following the methodology of [31], we categorized light rays intersecting the disk once, twice, and three or more times as direct, lensed, and photon ring trajectories, respectively. We then presented the photon trajectories in Fig. 4 and orbit numbers in Figs. 5 and 6 for both semi-disks at fixed inclination $\omega$ and polar angle $\eta$. The results reveal striking differences between the two semi-disks: the lensed emission region is larger on the co-side, and its size increases markedly with inclination, while the corresponding region on the counter-side shrinks under the same conditions.
@@
To establish the relationship between the accretion disk and points on the observer’s plane, we introduced transfer functions $r_{\rm m}(b)$ in Eq. (12) and $R_{\rm m}(b)$ in Eq. (13) for the counter-side and co-side semi-disks, respectively. These functions characterize the disk intersection positions of light rays received at each point on the observer’s plane. We presented the variation of these transfer functions with inclination angle $\omega$ and polar angle $\eta$ in Figs. 7 and 8. The first ($m=1$) transfer functions for both semi-disks exhibit progressively steeper slopes with increasing $\omega$ and $\eta$, accompanied by a reduction in the corresponding impact parameter ranges. In contrast, the second ($m=2$) and third ($m=3$) transfer functions of both semi-disks exhibit markedly distinct behaviors with increasing $\omega$ and $\eta$. Notably, when both the $\omega$ and $\eta$ approach $\pi/2$, the lensed emission region of the co-side semi-disk dominates the image contribution.
@@
Subsequently, by specifying the emission profiles of the thin disk models and incorporating the transfer functions, we calculated the received intensity at different polar angles $\eta$ for a fixed inclination $\omega$, with partial results presented in Fig. 10 and 11. Our analysis demonstrates that the co-side semi-disk exhibits more complex peak structures in the received intensity across varying polar angles $\eta$, whereas the counter-side semi-disk shows comparatively minor intensity variations. Moreover, we calculated the received intensity for each model at various $\omega$ and projected these results onto a two-dimensional plane to obtain the optical appearances of the three thin disk models under different inclinations, as shown in Fig. 12.
@@
Finally, we extended our discussion to optically thick versions of the three models, comparing their images with the thin disk cases using $\omega=17\pi/36$ as a representative example in Fig. 13. The results demonstrate that the thick disk’s obscuration effect eliminates certain bright rings and reduces the overall intensity by approximately $20\%$$40\%$ compared to thin disks.
@@ Appendix A Received intensities of Model-II and Model-III
In this appendix, we briefly present the received intensity profiles for selected polar angles $\eta$ in optically thin accretion disk Models II and III, shown in Figs. 1417, at inclination angles of $\pi/4$ and $3617\pi/$ 36.